OpenStudy (loser66):

What is 0^0 ? is it 1 or 0? Please, explain

OpenStudy (paxpolaris):

1

OpenStudy (loser66):

Explain, please.

OpenStudy (loser66):

(anything)^0 =1, you count exponent but 0^(anything) =0, you count the base Why do you choose the first one but the second one?

OpenStudy (paxpolaris):

by definition a^0 = 1 for any a... including 0

myininaya (myininaya):

\[\lim_{x \rightarrow 0}e^{\ln(x^x)}=e^{\lim_{x \rightarrow 0}x^x}=e^{\lim_{x \rightarrow 0} \frac{\ln(x)}{\frac{1}{x}}}=e^{\lim_{x \rightarrow 0}\frac{\frac{1}{x}}{-\frac{1}{x^2}}} \\ =e^{\lim_{x \rightarrow 0}\frac{x}{-1}}=e^{0}=1\] but we aren't talking about limits i always under 0^0 and 0/0 is indeterminate form it cannot be determined at 0 or 1

myininaya (myininaya):

it cannot be determined as 0 or 1

OpenStudy (paxpolaris):

enter 0^0 in calculator and you get 1. ... it is defined to be 1 to keep it consistent with all other a^0

myininaya (myininaya):

\[0^0=0^{n-n}=\frac{0^n}{0^n}=(\frac{0}{0})^n\]

OpenStudy (loser66):

I watch this video, at 17:00, the prof use 0^0 =1, I wonder why, That's why I make question here http://www.youtube.com/watch?v=KJuSx1EXdd8

myininaya (myininaya):

i always tell my students a^0=1 as long as a doesn't equal 0

ganeshie8 (ganeshie8):

\[\large (1-1)^n = \sum\limits_{k=0}^n \binom{n}{k}(-1)^{n-k}\] plugin n = 0

myininaya (myininaya):

it looks like that video is talking about limits

OpenStudy (loser66):

So, 2 profs, you and the prof in the video tap, teach the students (I am one of them) in 2 different ways? @myininaya . You guys confuse students . hahah

OpenStudy (loser66):

*tape

myininaya (myininaya):

i also don't see where you are talking about in the video where he says 0^0 is 1

OpenStudy (loser66):

At 15:00, he is talking about geometric sequence with r^n , at 17:00 , he expands the sequence as 1,0,0,0...

OpenStudy (loser66):

when r =0

OpenStudy (paxpolaris):

a^3 = a*a*a*1 a^2 = a*a*1 a^1 = a*1 a^0 = 1 this work for a=0 as well. @myininaya we know 0*3 = 0 even though 0/0 is not 3 ... same way we can define 0^0=1 even though we can't do 0^n/0^n

myininaya (myininaya):

it doesn't matter what he called 0^0 the sequence numbers still get closer to 0

myininaya (myininaya):

i will never believe 0^0 is 1

OpenStudy (paxpolaris):

BUT IT IS :'(

myininaya (myininaya):

\[0^0=0^{n-n}=\frac{0^n}{0^n}=(\frac{0}{0})^n \] how do you argue this then @PaxPolaris ?

OpenStudy (xapproachesinfinity):

eh no 0^0 is problematic so mathematicians tend to avoid it

myininaya (myininaya):

but i guess we could play with that arguement 0^5=0 but 0^5=0^{5-0}=0^5/0^0

myininaya (myininaya):

but still 0^0 is an indeterminate form this means it cannot be determined

OpenStudy (xapproachesinfinity):

what ever you do you will different results therefore we don't know what does that mean

OpenStudy (paxpolaris):

\[0^0=0^{n-n} \ne\frac{0^n}{0^n}=(\frac{0}{0})^n\]

myininaya (myininaya):

i know i already found a countexample for that arugment but there are other arguments against 0^0 being 1

ganeshie8 (ganeshie8):

If you want to believe in binomial theorem, define it as 1. You could define it as 0 and simply exuse binomial theorem and million other things your defintion contradicts with... below is not affected by whatever you choose \[\large \lim\limits_{x,y \to 0} x^y = DNE\]

OpenStudy (loser66):

Conclusion: Depend on what the current professor says, we define whether 0^0 =1 or 0^0 =0 LOL

ganeshie8 (ganeshie8):

\(\large \dfrac{0}{0}\) is undefinable based on my understanding we can define 0^0, 1^infty etc based on context..

OpenStudy (loser66):

Thanks you all. I really don't know this is a controversy problem in Math

OpenStudy (xapproachesinfinity):

if we believe that 0^0=1 a lot of things will get complicated lol so 0^0 is problematic @myininaya gave already some counterexamples

myininaya (myininaya):

well the counterexample i gave i was actually able to give a counterexample to that counterexample saying that logic probably wasn't the best route to choose -- bottom story is 0^0 cannot be determined (it is why they call it an indeterminate form)

myininaya (myininaya):

if 0^0 was a number you would never have needed to use l'hosptal rule for that form

ganeshie8 (ganeshie8):

Exactly! limit is a different story

ganeshie8 (ganeshie8):

basically in number theory and sequences/series you can assume 0^0 as 1. but in multivariable calculus or when you consider limits of two different functions, you need to consider it as indeterminate. be aware of whats going on, then you will be fine..

OpenStudy (paxpolaris):

much ado about nothing (about 0)

myininaya (myininaya):

And I do agree with @ganeshie8 on the part you can define 0^0 as 1 for certain purposes.